A company’s stock is currently selling for $125 per share. The company has just paid a...

  1. A company’s stock is currently selling for $125 per share. The company has just paid a dividend of $4.25 per share. Investors expect this dividend to grow at the rate of 6% per year for the next 7 years and at a rate of 4% per year thereafter.

  1. Calculate the cost of equity for this company.
  2. Calculate, using either Goal Seek or Solver, a current stock price that would yield a cost of equity of 9%.
Answer #1

Where 1 st year end dividend is = 4.25*1.06= 4.505.

so cost of capital is 7.99%.From cell E2.

b)

Calculation of price of stock at 9% cost of capital using goal seek formula.

using goal seek formula, Calculating cost of stock by changing E2 cell to 1.09. and D11 is the stock price
Year Devidend Pv factor Devidend*PV factor 1.09
1 4.505 =1/E2 =B3*C3
2 =B3*1.06 =C3/E2 =B4*C4
3 =B4*1.06 =C4/E2 =B5*C5
4 =B5*1.06 =C5/E2 =B6*C6
5 =B6*1.06 =C6/E2 =B7*C7
6 =B7*1.06 =C7/E2 =B8*C8
7 =B8*1.06 =C8/E2 =B9*C9
7 =B9*(1.04)/(E2-1-0.04) =C8/E2 =B10*C10
=SUM(D3:D10)
using goal seek formula, Calculating cost of stock by changing E2 cell to 1.09. and D11 is the stock price
Year Devidend Pv factor Devidend*PV factor 1.09
1 4.505 0.917431 4.133027523
2 4.7753 0.84168 4.019274472
3 5.061818 0.772183 3.908652239
4 5.365527 0.708425 3.801074654
5 5.687459 0.649931 3.696457921
6 6.028706 0.596267 3.594720547
7 6.390429 0.547034 3.495783284
7 132.9209 0.547034 72.7122923
99.36128294

So stock price at 9 % is 99.36.